OpenStudy (anonymous):

\[\int\limits_{?}^{?}(x+1/x)^2dx\]

OpenStudy (wasiqss):

u can open the square , and then do long division

OpenStudy (anonymous):

If you open the square, you have \[\int\limits_{}^{}x^2 +2 +1/x^2+k\] I just forgot how to integrate...

OpenStudy (mimi_x3):

is it: \[\int\limits\left(\frac{x+1}{x}\right) ^{2} \]

OpenStudy (wasiqss):

u know wut long divison is

OpenStudy (anonymous):

No.. \[\int\limits_{}^{}[(x) + (1/x)]^2\]

OpenStudy (wasiqss):

no when u simplify it wud b {(1+(1/x)}^2

OpenStudy (anonymous):

x^3/3 + 2x - 1/x +c

OpenStudy (wasiqss):

i hate sumone pops out from middle jus telling answer, i was making her understand!

OpenStudy (anonymous):

just convert (x+1/x)^2 to x^2+ 1/x^2 +2 and use the integration rules

OpenStudy (anonymous):

Why is the second to the last one -1/x?

OpenStudy (wasiqss):

arnab she doesnt know long division!

OpenStudy (anonymous):

because 1/x^2 = x^-2 when u integrate x^n, it will be x^(n+1)/(n+1)

OpenStudy (anonymous):

@ order, u dont know the formula of (a+b)^2??

OpenStudy (anonymous):

Um I think so... how does x^-2 become -1x?

OpenStudy (anonymous):

I mean -1/x

OpenStudy (anonymous):

integrate x^-2 so, it will be x^-1/(-1)=-1/x

OpenStudy (mimi_x3):

\[\int\limits \left(x+\frac{1}{x}\right) ^{2} = \int\limits x^{2}+\frac{1}{x^{2}} +2 \] \[ \frac{x^{3}}{3} +x^-2 + 2x \] \[\frac{x^{3}}{3} +\frac{x^{-1}}{-1} +2x + C\]

OpenStudy (anonymous):

Ok, that makes a lot more sense :) Thanks Mimi, and also you Arnab... I just couldn't understand when you wrote it...

OpenStudy (mimi_x3):

woops typo: \[ \frac{x^{3}}{3} +x^{-2} + 2x \]

OpenStudy (anonymous):

welcome :)